# Riemannian Metric Learning: Closer to You than You Imagine

### v1

@@ Acknowledgments
One can approach this problem via optimal transport [152]: For any $j\in[T-1]$, a transport map $T^{j}$ is learned such that $P^{j+1}=T^{j}\#P^{j}$ is the pushforward measure of $P^{j}$ by $T^{j}$. When using the Euclidean metric as base metric, the interpolation becomes $P^{j+t}=(tT^{j}+(1-t)\mathrm{Id})\#P^{j}$ for any $t\in[0,1]$. However, this Euclidean interpolation can put some mass in regions of the space which are unrealistic. A conceptual illustration is given in Figure 8. That is why Scarvelis and Solomon [157] and Huguet et al. [95] proposed to learn a Riemannian metric to be used in the optimal transport problem. In the end, for any $x$ in the observational space, defining $\gamma(x,t)$ as the geodesics between $x$ and $T^{j}(x)$ for any $j\in[T-1]$, the interpolation becomes $P^{j+t}=\gamma(\cdot,t)\#P^{j}$. The method improved the results on a bird migration dataset [^11] https://birdsoftheworld.org/bow/species/snogoo/cur/movement and an scRNA dataset [158].
Flow matching [117] is another option, where the interpolation of distributions is achieved by integrating a flow equation $\frac{\mathop{}\!\mathrm{d}X}{\mathop{}\!\mathrm{d}t}=v_{t}(X)$. Here, the velocity field is learned in order to match a quantity related to an interpolating function $I\colon\mathcal{M}^{2}\times[0,1]\to\mathcal{M}$. This function is a prior on the shape of trajectory we want after integrating the flow equation. Most of the time, $I$ is chosen to be linear because of its simplicity. However, as for Euclidean optimal transport interpolation, this choice is not optimal since the interpolation might leave the data manifold as depicted in Figure 8. Again, a Riemannian metric can be learned from the dataset and geodesics can be used to define the desired interpolant as $I(x,y,t)\approx\mathrm{Exp}_{x}(t\mathrm{Log}_{x}(y))$ using neural network approximations [100, 170]. Kapusniak et al. [100] and Sun et al. [170] successfully used Riemannian metric learning within flow matching for single-cell trajectory inference on the CITE-seq and Multiome datasets from a NeurIPS competition [29], where both methods greatly improved on the state-of-the-art.
@@
More generally, Riemannian metric learning finds also applications in sampling [44, 34, 35] and generative modeling [5, 63, 8]. In these applications, the metric $\mathfrak{g}$ and its volume $\text{vol}_{\mathfrak{g}}$ are particularly involved. In [68], the authors demonstrate that sampling algorithms such as Hamiltonian Monte Carlo (HMC) [24] can be improved by using a Riemannian metric related to the distribution. Chadebec et al. [34] built on this idea by learning the metric on a latent space in a variational autoencoders to perform data augmentation for brain MRI [35]. Learning more structure in the latent space is particularly interesting [35, 180] when datasets are small and high dimensional—such as MRI datasets. Sun et al. [170] suggest to sample uniformly according to the volume $\text{vol}_{\mathfrak{g}}$ in the Riemannian latent space in order to avoid selection bias due to observations using the unadajusted Langevin algorithm as depicted on the left of Figure 9. Cui et al. [44] remind us that the convergence guarantee of Langevin dynamics to their invariant distribution is related to Poincaré inequalities. Thus, they optimize the Poincaré constant to accelerate convergence by learning a Riemannian metric on the state space. Therefore, Riemannian metric learning can be a tool to improve sampling efficiency motivated by convergence results.
In an unsupervised learning context, Arvanitidis et al. [5] suggested to learn a Gaussian mixture on a Riemannian manifold adapted to a sleep stages dataset from Physionet. They adapted the geometry of the Gaussian in order to avoid holes in the data distribution. Using autoencoders and variational autoencoders, several papers [6, 63, 9, 8, 47] have underlined that learning the Riemannian structure of the latent space improves uncertainty estimation and interpretability.
For instance, de Kruiff et al. [47] they learn an isometric map between the latent space and the observational space in order to preserve distance related to biological properties of observed proteins. By using the Riemannian structure of their latent manifold, they can generate new proteins with specific biological properties with simple geodesic interpolation [47, Figure 5] as depicted on the right of Figure 9. A score-based variant of the method is given in [49] and its theoretical underpinnings are discussed in more detail in [48].
Figure 9: On the left, a Riemannian manifold strcuture is learned [170] (licensed under CC BY 4.0) using embryoid data and data augmentation is performed by sampling according to the metric volume. On the right, synthetic proteins are generated in [47] (licensed under CC BY 4.0) by interpolating a protein with high hydrophobic moment and low charge $x$ and an another one with a low hydrophobic moment and high charge $y$.
@@
Furthermore, domain adaptation can also benefit from Riemannian metric learning using parallel transport. For instance, in cardiac motion modeling, when a motion is estimated as a geodesic passing through $q_{0}$ in the source domain, we can transport this trajectory to a point $q_{1}$ in the target domain by transporting its velocity in parallel along the geodesic going from $q_{0}$ to $q_{1}$ [73]. This is illustrated on the right of Figure 10 with bear shapes instead of cardiac motions for the sake of simplicity.
Tam et al. [172] apply a similar reasoning when working on a Riemannian space of probability distribution. Points in the source and target domains can be represented using normal coordinates at their Karcher means $p_{S}$ and $p_{T}$ (3.2), respectively. Then the normal representation of the target features can be transported to the source domain using the parallel transport following a trajectory going from $p_{T}$ to $p_{S}$. This is illustrated on the left of Figure 10. In practice, to reduce the computational burden, Tam et al. [172] learn a statistical model (linear regression, PCA) on the normal representations in the source domain, then they transport the statistical model to the target domain. Crucially, this transport is independant of the number of data points [172].
@@
Now that we have introduced a range of applications for metric learning, we will turn back to its technical aspects, i.e. how to make the abstract problem from Section 3.1 more concrete and how to solve it. While the choice of the objective function $\mathfrak{L}$ is at the crux of the problem (Section 4.2), the choice of the parametrization (Section 4.1) will deeply influence its flexibility and its computational tractability. Interestingly, the optimization methods related to Riemannian metric learning (Section 4.3) are less specific to the problem than one would expect, the discussion on this issue is continued in Section 5.
@@
The first aspect we discuss is how to parametrize the learned metric, or, in other words, which ansatz space $\mathcal{G}_{\theta}\subset\mathcal{G}_{\mathcal{M}}$ to select for the learning problem (3.1). As highlighted in Section 3.1, the manifold of metrics $\mathcal{G}_{\mathcal{M}}$ on a manifold $\mathcal{M}$ is infinite-dimensional. Thus, to perform optimization, we must restrict our focus to a finite-dimensional subspace. Over the years, various approaches have been proposed for this purpose, and we will summarize and critically discuss the most significant ones.
@@
The idea of a constant metric has also been extended to other types of manifolds. A common extension is to matrix manifolds, where the Frobenius inner product can be weighted, for example, via expressions like $\mathrm{tr}(V^{T}GW)$ or $\mathrm{tr}(GV^{T}W)$, or by adapting the distance functions introduced in Section 2.1. Such approaches have been applied to SPD matrix manifolds [94, 41, 183, 208] and Grassmannian manifolds [75, 78, 208].
@@
An alternative method involves discretizing the manifold with a (simplicial) grid and assigning a constant metric to each grid cell. This strategy, inspired by finite element methods, was explored by Cui et al. [44] and Qiu and Dai [148], who applied it from a computational perspective. However, discretizing high dimensional manifolds using a grid quickly yields computational challenges.
@@
Another neural network-based approach takes inspiration from the Hamiltonian formulation of Riemannian geometry (see Section 2) by directly parametrizing the Hamiltonian, as explored in [176, 40]. This perspective allows for learning the underlying geometric structure through dynamics.
Finally, van Veldhuizen [180], Chadebec et al. [34, 35], Scarvelis and Solomon [157] combine neural networks with interpolation techniques and variational autoencoder latent distribution. In their approach, the metric is first parametrized at key points using a neural network of the form $G_{i}=A_{i}^{\vphantom{T}}A_{i}^{T}+\eta\mathrm{I}_{d}$, and then extended across the manifold via Gaussian kernel interpolation.
@@
In many cases, the primary objective is to ensure that geodesics remain close to an observational dataset $\mathcal{D}=(x_{i})_{i}$, as real-world observations often lie on an unknown manifold ~$\widetilde{\mathcal{M}}\subset\mathbb{R}^{d}$. To achieve this, Arvanitidis et al. [5, 6] and Kapusniak et al. [100] suggest defining a metric $\mathfrak{g}$ on $\mathbb{R}^{d}$ as
@@
for any point $x\in\mathbb{R}^{d}$ with coordinates $x=(x^{j})_{j\in[d]}$ where $\epsilon>0$ and $\sigma>0$ are hyperparameters to be tuned. With this choice, when $x$ is far from the dataset $\mathcal{D}$, $h(x)$ will be small because of the exponential term contrary to the metric’s minimal eigenvalue since $\mathfrak{g}_{x}\approx\epsilon^{-1}\mathrm{I}_{d}$. Thus, when $x$ is far from the dataset $\mathcal{D}$, it induces large energies of displacement as the sketched in Figure 3. To increase the contrast, the hyperparameters are chosen such that $h(x)\approx 1$ when $x$ is in the dataset. This method has been shown to be effective in tasks such as flow matching, where geodesics play a critical role [100].
@@
where $\mathop{}\!\mathrm{d}_{x}f$ denotes the differential of $f$ at point $x$. This pullback metric is valid as long as $f$ is locally injective. The map $f$ can take various forms depending on the application. In the simplest case, $f$ is a linear map, which brings us back to the scenario of Mahalanobis distance learning. More commonly, nonlinear forms of $f$ are used including kernel methods [92, 131, 175, 72] and neural networks [63, 99, 123, 206].
A frequent application of pullback metrics involves autoencoder architectures [200, 191, 171, 170, 142, 175, 99, 16] where the metric can be induced by either the encoder or the decoder depending on the task requirements. When $f\sim P$ is random or not unique in the context, relying on the expected value $\mathbb{E}_{f\sim P}(\mathop{}\!\mathrm{d}_{x}f)$ instead of $\mathop{}\!\mathrm{d}_{x}f$ is an interesting extension account for the probabilistic nature of the space geometry [175, 60, 171]. Arvanitidis et al. [9] designed the metric according to a learned prior in the latent space and not using the pullback to gain robustness while still having a similar structure.
While some method directly explicitly define the pullback metric [72, 175, 123, 99], others, such as [27, 95, 93, 77, 131, 102, 11, 92, 61, 52, 31, 39, 56, 82] focus on learning a local approximation of the distance using the map-induced distance
@@
Note that $\mathbf{d}_{f}$ is only a local approximation of the geodesic distance, as it does not account for the full geometry of the manifold as explained in Section 2.
@@
Such graph-based methods are commonly used to learn distances in a non-parametric manner, based on the assumption that the dataset is isometrically embedded in an ambient space. The nearest-neighbor graph is then computed leading to algorithms like Isomap [173] and its extensions such as [28].
@@
Once one has parametrized the Riemannian metric $\mathfrak{g}$, one has to compute the Riemannian objects of interest $\mathcal{F}(\mathfrak{g})=(\mathbf{d}_{\mathfrak{g}},\mathrm{Exp}_{\mathfrak{g}},\mathfrak{T}_{\mathfrak{g}},\text{vol}_{\mathfrak{g}})$ as explained in Section 3. Here, we provide a few elaborations on how to do this.
@@
For nonlinear explicit forms, no universally accepted gold standard currently exists. The metric volume $\text{vol}_{\mathfrak{g}}$ requires computing a determinant at each point $x\in\mathcal{M}$. The exponential map $\mathrm{Exp}_{\mathfrak{g}}$ and the parallel transport $\mathfrak{T}_{\mathfrak{g}}$ can be approximated by numerically integrating the corresponding second-order differential equations, see Section 2.
@@
    where $\tilde{\mathbf{d}}_{\mathfrak{g}}(x_{k},x_{k+1})^{2}$ is a local approximation of the squared distance, e.g. $\tilde{\mathbf{d}}_{\mathfrak{g}}(x_{k},x_{k+1})^{2}\approx\mathfrak{g}_{x_{k}}(x_{k+1}-x_{k},x_{k+1}-x_{k})$. Here, $N$ must be sufficiently large to ensure a realistic solution [83, 7] and the convergence of this approach for $N$ going to infinity was investigated by Rumpf and Wirth [150]. The minimization can be performed in various ways. A recent fast and robust approach suggests using Gaussian processes [7] combined with fixed-point iterations to avoid the computational overhead of Jacobian processing. While such methods may be less precise than the shooting method, in some cases shooting methods are not feasible especially in high dimensions.
@@
The first set of objectives based on classification we consider are usually referred to as contrastive losses [37, 38, 46, 61, 70, 74, 75, 78, 77, 80, 92, 94, 93, 102, 189, 191, 196, 52]. Here, one assume two sets as input: a set $S_{p}\subset\mathcal{X}\times\mathcal{X}$ of examples that should be classified as similar (e.g. have the same label) and another set $S_{n}\subset\mathcal{X}\times\mathcal{X}$ of examples that should be classified as different. These sets, especially the latter, are typically assumed to contain “challenging” examples for a classification algorithm, i.e. such examples that are close to an envisioned decision boundary (in the Euclidean norm). Hence, the objective in metric learning is to achieve a clearer separation of these examples, i.e. to maximize the distance for elements of $S_{n}$, while keeping similar points close, i.e. to minimize the distance for elements of $S_{p}$. Various different objectives have been proposed that formalize this idea, for example, as
@@
If one has a notion of density of the samples associated with them, i.e. some form of weights, these can of course also be integrated into the objective. This approach to regress distances has, for example, been investigated in [47, 48, 49, 63, 148].
@@
To this end, Louis et al. [123] consider an autoencoder setup where the metric on the data manifold is given by the push-forward of the Euclidean metric on the latent space (cf. Section 4.1). This means geodesics on the data manifold are given by the images of linear trajectories in the latent space, which are much more tractable to work with. Furthermore, they assume that all observed trajectories are parallel to each other (and a coordinate axis), which means that each trajectory is described by its constant speed, its starting time relative to the observations, and its offset to the coordinate axis. These quantities are then optimized together with the parameters of the autoencoder determining the Riemannian metric considering a objective of the form (4.15) and an additional regularization on them. This idea has been extended by Sauty and Durrleman [155] to variational autoencoders also modeling the data distribution (see above). Here, the regression objective is obtained by comparing points in the latent space whereas Louis et al. [123] compare points in the original data space after decoding.
This idea can be generalized to nonlinear $\exp$-parallel curves [159] (see Section 3.2.2), which are not necessarily geodesics. Sauty and Durrleman [156] employ this idea while directly learning a metric on the data manifold. Thus, they approximate the geodesic and the parallel curves using a Hamiltonian shooting approach. In contrast, Gruffaz et al. [72] obtain the metric using a diffeomorphic deformation of a reference manifold (usually the Euclidean space) with known closed forms for the geodesic and parallel curves.
@@
As explained in Section 2.1, one can learn a metric on the space of probability measures on a manifold $\mathcal{M}$ by learning a Riemannian metric on this base manifold. Hence, a range of works as focused on objectives for Riemannian metric learning that incorporates optimal transport based on it.
Considering the classification of histograms, Cuturi and Avis [45] adapt their objective from the contrastive losses in distance metric learning and arrive at the loss $\mathfrak{L}(\mathfrak{g})=\sum_{(\mu,\nu)\in S_{p}}W_{2}(\mu,\nu)-\sum_{(\mu,\nu)\in S_{n}}W_{2}(\mu,\nu)$, where $W_{2}$ is the Wasserstein distance induced by $\mathfrak{g}$ as defined in (2.25). In contrast, Zhou et al. [206] minimize the Wasserstein distance between subsets of data points (representing different source domains in a domain adaption setup) while combining it with a triplet loss (see Section 4.2.1) over all data points. Likewise, Kerdoncuff et al. [103] optimize the Wasserstein distance between two sets of data points, one with labels and one without, with an additional penalty on the transport plan preventing two differently labeled points being sent to the same target. They also combine this with classification-based losses on the points with known labels.
@@
Thus, in the absence of closed form solutions, most works turn towards iterative descent algorithms to solve the optimization problem. Here, the most fundamental approach is to apply gradient descent or its stochastic extension which was used in [14, 38, 70, 71, 103, 206, 161, 102, 195, 205, 139]. To speed-up the convergence of the descent, some works, such as [8, 34, 35, 100, 121, 176, 155, 156, 154, 157], use acceleration techniques such as Nesterov’s acceleration or the Adam algorithm. Others [84, 163] use L-BFGS to approximate the Hessian of the objective and achieve faster convergence.
These iterative methods typically face two challenges: (i) computing the gradient, and (ii) ensuring that the learned metric is positive definite for explicit parametrizations. The first challenge traditionally required manual computation of the gradient of the objective which has been mostly replaced by the usage of automatic differentiation frameworks in more recent approaches. Furthermore, the objective might not admit a classical gradient, e.g. if it involves solving an optimization problem as in the case of optimal transport-based objectives. Thus, some works explicitly use subgradient methods [45, 102, 194, 139]. The second challenge is often solved via the parametrization of the metric (see Section 4.1) and thus mostly comes up in the optimization for constant metrics. In these cases, it is typically [70, 195] solved by projected gradient descent, i.e. alternating a gradient step with projecting the metric matrix onto cone of positive definite matrices.

### v2

@@ Acknowledgments
One can approach this problem via optimal transport [157]: For any $j\in[T-1]$, a transport map $T^{j}$ is learned such that $P^{j+1}=T^{j}\#P^{j}$ is the pushforward measure of $P^{j}$ by $T^{j}$. When using the Euclidean metric as base metric, the interpolation becomes $P^{j+t}=(tT^{j}+(1-t)\mathrm{Id})\#P^{j}$ for any $t\in[0,1]$. However, this Euclidean interpolation can put some mass in regions of the space which are unrealistic. A conceptual illustration is given in Figure˜8. That is why Scarvelis and Solomon [162] and Huguet et al. [98] proposed to learn a Riemannian metric to be used in the optimal transport problem. In the end, for any $x$ in the observational space, defining $\gamma(x,t)$ as the geodesics between $x$ and $T^{j}(x)$ for any $j\in[T-1]$, the interpolation becomes $P^{j+t}=\gamma(\cdot,t)\#P^{j}$. The method improved the results on a bird migration dataset [^11] https://birdsoftheworld.org/bow/species/snogoo/cur/movement and an scRNA dataset [163].
Flow matching [120] is another option, where the interpolation of distributions is achieved by integrating a flow equation $\frac{\mathop{}\!\mathrm{d}X}{\mathop{}\!\mathrm{d}t}=v_{t}(X)$. Here, the velocity field is learned in order to match a quantity related to an interpolating function $I\colon\mathcal{M}^{2}\times[0,1]\to\mathcal{M}$. This function is a prior on the shape of trajectory we want after integrating the flow equation. Most of the time, $I$ is chosen to be linear because of its simplicity. However, as for Euclidean optimal transport interpolation, this choice is not optimal since the interpolation might leave the data manifold as depicted in Figure˜8. Again, a Riemannian metric can be learned from the dataset and geodesics can be used to define the desired interpolant as $I(x,y,t)\approx\mathrm{Exp}_{x}(t\mathrm{Log}_{x}(y))$ using neural network approximations [103, 175]. Kapusniak et al. [103] and Sun et al. [175] successfully used Riemannian metric learning within flow matching for single-cell trajectory inference on the CITE-seq and Multiome datasets from a NeurIPS competition [32], where both methods greatly improved on the state-of-the-art.
@@
More generally, Riemannian metric learning finds also applications in sampling [47, 37, 38] and generative modeling [6, 66, 9]. In these applications, the metric $\mathfrak{g}$} and its volume $\text{vol}_{\mathfrak{g}}$ are particularly involved. In [71], the authors demonstrate that sampling algorithms such as Hamiltonian Monte Carlo (HMC) [26] can be improved by using a Riemannian metric related to the distribution. Chadebec et al. [37] built on this idea by learning the metric on a latent space in a variational autoencoders to perform data augmentation for brain MRI [38]. Learning more structure in the latent space is particularly interesting [38, 187] when datasets are small and high dimensional—such as MRI datasets. Sun et al. [175] suggest to sample uniformly according to the volume $\text{vol}_{\mathfrak{g}}$ in the Riemannian latent space in order to avoid selection bias due to observations using the unadajusted Langevin algorithm as depicted on the left of Figure˜9. Cui et al. [47] remind us that the convergence guarantee of Langevin dynamics to their invariant distribution is related to Poincaré inequalities. Thus, they optimize the Poincaré constant to accelerate convergence by learning a Riemannian metric on the state space. Therefore, Riemannian metric learning can be a tool to improve sampling efficiency motivated by convergence results.
In an unsupervised learning context, Arvanitidis et al. [6] suggested to learn a Gaussian mixture on a Riemannian manifold adapted to a sleep stages dataset from Physionet. They adapted the geometry of the Gaussian in order to avoid holes in the data distribution. Using autoencoders and variational autoencoders, several papers [7, 66, 10, 9, 50, 154] have underlined that learning the Riemannian structure of the latent space improves uncertainty estimation and interpretability.
For instance, de Kruiff et al. [50] they learn an isometric map between the latent space and the observational space in order to preserve distance related to biological properties of observed proteins. By using the Riemannian structure of their latent manifold, they can generate new proteins with specific biological properties with simple geodesic interpolation [50, Figure 5] as depicted on the right of Figure˜9. A score-based variant of the method is given in [52] and its theoretical underpinnings are discussed in more detail in [51].
Figure 9: On the left, a Riemannian manifold strcuture is learned [175] (licensed under $C$ BY 4.0) using embryoid data and data augmentation is performed by sampling according to the metric volume. On the right, synthetic proteins are generated in [50] (licensed under $C$ BY 4.0) by interpolating a protein with high hydrophobic moment and low charge $x$ and an another one with a low hydrophobic moment and high charge $y$.
@@
Furthermore, domain adaptation can also benefit from Riemannian metric learning using parallel transport. For instance, in cardiac motion modeling, when a motion is estimated as a geodesic passing through $q_{0}$ in the source domain, we can transport this trajectory to a point $q_{1}$ in the target domain by transporting its velocity in parallel along the geodesic going from $q_{0}$ to $q_{1}$ [76]. This is illustrated on the right of Figure˜10 with bear shapes instead of cardiac motions for the sake of simplicity.
Tam et al. [179] apply a similar reasoning when working on a Riemannian space of probability distribution. Points in the source and target domains can be represented using normal coordinates at their Karcher means $p_{S}$ and $p_{T}$ (3.2), respectively. Then the normal representation of the target features can be transported to the source domain using the parallel transport following a trajectory going from $p_{T}$ to $p_{S}$. This is illustrated on the left of Figure˜10. In practice, to reduce the computational burden, Tam et al. [179] learn a statistical model (linear regression, PCA) on the normal representations in the source domain, then they transport the statistical model to the target domain. Crucially, this transport is independant of the number of data points [179].
@@
Now that we have introduced a range of applications for metric learning, we will turn back to its technical aspects, i.e. how to make the abstract problem from Section˜3.1 more concrete and how to solve it. While the choice of the objective function $\mathfrak{L}$ is at the crux of the problem (Section˜4.2), the choice of the parametrization (Section˜4.1) will deeply influence its flexibility and its computational tractability. Interestingly, the optimization methods related to Riemannian metric learning (Section˜4.3) are less specific to the problem than one would expect, the discussion on this issue is continued in Section˜5.
@@
The first aspect we discuss is how to parametrize the learned metric, or, in other words, which ansatz space $\mathcal{G}_{\theta}\subset\mathcal{G}_{\mathcal{M}}$ to select for the learning problem (3.1). As highlighted in Section˜3.1, the manifold of metrics $\mathcal{G}_{\mathcal{M}}$ on a manifold $\mathcal{M}$ is infinite-dimensional. Thus, to perform optimization, we must restrict our focus to a finite-dimensional subspace. Over the years, various approaches have been proposed for this purpose, and we will summarize and critically discuss the most significant ones.
@@
The idea of a constant metric has also been extended to other types of manifolds. A common extension is to matrix manifolds, where the Frobenius inner product can be weighted, for example, via expressions like $\mathrm{tr}(V^{T}GW)$ or $\mathrm{tr}(GV^{T}W)$, or by adapting the distance functions introduced in Section˜2.1. Such approaches have been applied to SPD matrix manifolds [97, 44, 190, 215] and Grassmannian manifolds [78, 81, 215].
@@
An alternative method involves discretizing the manifold with a (simplicial) grid and assigning a constant metric to each grid cell. This strategy, inspired by finite element methods, was explored by Cui et al. [47] and Qiu et al. [152], who applied it from a computational perspective. However, discretizing high dimensional manifolds using a grid quickly yields computational challenges.
@@
Another neural network-based approach takes inspiration from the Hamiltonian formulation of Riemannian geometry (see Section˜2) by directly parametrizing the Hamiltonian, as explored in [183, 43]. This perspective allows for learning the underlying geometric structure through dynamics.
Finally, Chadebec et al. [37, 38], Scarvelis and Solomon [162], and van Veldhuizen [187] combine neural networks with interpolation techniques and variational autoencoder latent distribution. In their approach, the metric is first parametrized at key points using a neural network of the form $G_{i}=A_{i}^{\vphantom{T}}A_{i}^{T}+\eta\mathrm{I}_{d}$, and then extended across the manifold via Gaussian kernel interpolation.
@@
In many cases, the primary objective is to ensure that geodesics remain close to an observational dataset $\mathcal{D}=(x_{i})_{i}$, as real-world observations often lie on an unknown manifold $\widetilde{\mathcal{M}}\subset\mathbb{R}^{d}$. To achieve this, Arvanitidis et al. [6, 7] and Kapusniak et al. [103] suggest defining a metric $\mathfrak{g}$} on $\mathbb{R}^{d}$ as
@@
for any point $x\in\mathbb{R}^{d}$ with coordinates $x=(x^{j})_{j\in[d]}$ where $\epsilon>0$ and $\sigma>0$ are hyperparameters to be tuned. With this choice, when $x$ is far from the dataset $\mathcal{D}$, $h(x)$ will be small because of the exponential term contrary to the metric’s minimal eigenvalue since ${\mathfrak{g}}_{x}\approx\epsilon^{-1}\mathrm{I}_{d}$. Thus, when $x$ is far from the dataset $\mathcal{D}$, it induces large energies of displacement as the sketched in Figure˜3. To increase the contrast, the hyperparameters are chosen such that $h(x)\approx 1$ when $x$ is in the dataset. This method has been shown to be effective in tasks such as flow matching, where geodesics play a critical role [103].
@@
where $\mathop{}\!\mathrm{d}_{x}f$ denotes the differential of $f$ at point $x$. This pullback metric is valid as long as $f$ is locally injective. The map $f$ can take various forms depending on the application. In the simplest case, $f$ is a linear map, which brings us back to the scenario of Mahalanobis distance learning. More commonly, nonlinear forms of $f$ are used including kernel methods [95, 134, 182, 75] and neural networks [66, 102, 126, 213, 29].
A frequent application of pullback metrics involves autoencoder architectures [207, 198, 178, 175, 146, 182, 102, 18, 1, 29] where the metric can be induced by either the encoder or the decoder depending on the task requirements. When $f\sim P$ is random or not unique in the context, relying on the expected value $\mathbb{E}_{f\sim P}(\mathop{}\!\mathrm{d}_{x}f)$ instead of $\mathop{}\!\mathrm{d}_{x}f$ is an interesting extension account for the probabilistic nature of the space geometry [182, 63, 178, 154]. Arvanitidis et al. [10] designed the metric according to a learned prior in the latent space and not using the pullback to gain robustness while still having a similar structure.
While some method directly explicitly define the pullback metric [75, 182, 126, 102, 29], others, such as [30, 98, 96, 80, 134, 105, 12, 95, 64, 55, 34, 42, 59, 85] focus on learning a local approximation of the distance using the map-induced distance
@@
Note that $\mathbf{d}_{f}$ is only a local approximation of the geodesic distance, as it does not account for the full geometry of the manifold as explained in Section˜2.
@@
Such graph-based methods are often used to learn distances in a non-parametric manner, based on the assumption that the dataset is isometrically embedded in an ambient space. The nearest-neighbor graph is then computed leading to algorithms like Isomap [180] and its extensions such as [31].
@@
Once one has parametrized the Riemannian metric $\mathfrak{g}$}, one has to compute the Riemannian objects of interest $\mathcal{F}({\mathfrak{g}})=(\mathbf{d}_{\mathfrak{g}},\mathrm{Exp}_{\mathfrak{g}},\mathfrak{T}_{\mathfrak{g}},\text{vol}_{\mathfrak{g}})$ as explained in Section˜3. Here, we provide a few elaborations on how to do this.
@@
For nonlinear explicit forms, no universally accepted gold standard currently exists. The metric volume $\text{vol}_{\mathfrak{g}}$ requires computing a determinant at each point $x\in\mathcal{M}$. The exponential map $\mathrm{Exp}_{\mathfrak{g}}$ and the parallel transport $\mathfrak{T}_{\mathfrak{g}}$ can be approximated by numerically integrating the corresponding second-order differential equations, see Section˜2.
@@
    where $\tilde{\mathbf{d}}_{\mathfrak{g}}(x_{k},x_{k+1})^{2}$ is a local approximation of the squared distance, e.g.
    $\tilde{\mathbf{d}}_{\mathfrak{g}}(x_{k},x_{k+1})^{2}\approx{\mathfrak{g}}_{x_{k}}(x_{k+1}-x_{k},x_{k+1}-x_{k})$. Here, $N$ must be sufficiently large to ensure a realistic solution [86, 8] and the convergence of this approach for $N$ going to infinity was investigated by Rumpf and Wirth [155]. The minimization can be performed in various ways. A recent fast and robust approach suggests using Gaussian processes [8] combined with fixed-point iterations to avoid the computational overhead of Jacobian processing. While such methods may be less precise than the shooting method, in some cases shooting methods are not feasible especially in high dimensions.
@@
The first set of objectives based on classification we consider are usually referred to as contrastive losses [40, 41, 49, 64, 73, 77, 78, 81, 80, 83, 95, 97, 96, 105, 196, 198, 203, 55, 17]. Here, one assume two sets as input: a set $S_{p}\subset\mathcal{X}\times\mathcal{X}$ of examples that should be classified as similar (e.g. have the same label) and another set $S_{n}\subset\mathcal{X}\times\mathcal{X}$ of examples that should be classified as different. These sets, especially the latter, are typically assumed to contain “challenging” examples for a classification algorithm, i.e. such examples that are close to an envisioned decision boundary (in the Euclidean norm). Hence, the objective in metric learning is to achieve a clearer separation of these examples, i.e. to maximize the distance for elements of $S_{n}$, while keeping similar points close, i.e. to minimize the distance for elements of $S_{p}$. Various different objectives have been proposed that formalize this idea, for example, as
@@
If one has a notion of density of the samples associated with them, i.e. some form of weights, these can of course also be integrated into the objective. This approach to regress distances has, for example, been investigated in [50, 51, 52, 66, 152, 29].
@@
To this end, Louis et al. [126] consider an autoencoder setup where the metric on the data manifold is given by the push-forward of the Euclidean metric on the latent space (cf. Section˜4.1). This means geodesics on the data manifold are given by the images of linear trajectories in the latent space, which are much more tractable to work with. Furthermore, they assume that all observed trajectories are parallel to each other (and a coordinate axis), which means that each trajectory is described by its constant speed, its starting time relative to the observations, and its offset to the coordinate axis. These quantities are then optimized together with the parameters of the autoencoder determining the Riemannian metric considering a objective of the form (4.15) and an additional regularization on them. This idea has been extended by Sauty and Durrleman [160] to variational autoencoders also modeling the data distribution (see above). Here, the regression objective is obtained by comparing points in the latent space whereas Louis et al. [126] compare points in the original data space after decoding.
This idea can be generalized to nonlinear $\exp$-parallel curves [164] (see Section˜3.2.2), which are not necessarily geodesics. Sauty and Durrleman [161] employ this idea while directly learning a metric on the data manifold. Thus, they approximate the geodesic and the parallel curves using a Hamiltonian shooting approach. In contrast, Gruffaz et al. [75] obtain the metric using a diffeomorphic deformation of a reference manifold (usually the Euclidean space) with known closed forms for the geodesic and parallel curves.
@@
As explained in Section˜2.1, one can learn a metric on the space of probability measures on a manifold $\mathcal{M}$ by learning a Riemannian metric on this base manifold. Hence, a range of works as focused on objectives for Riemannian metric learning that incorporates optimal transport based on it.
Considering the classification of histograms, Cuturi and Avis [48] adapt their objective from the contrastive losses in distance metric learning and arrive at the loss $\mathfrak{L}({\mathfrak{g}})=\sum_{(\mu,\nu)\in S_{p}}W_{2}(\mu,\nu)-\sum_{(\mu,\nu)\in S_{n}}W_{2}(\mu,\nu)$, where $W_{2}$ is the Wasserstein distance induced by $\mathfrak{g}$} as defined in (2.25). In contrast, Zhou et al. [213] minimize the Wasserstein distance between subsets of data points (representing different source domains in a domain adaption setup) while combining it with a triplet loss (see Section˜4.2.1) over all data points. Likewise, Kerdoncuff et al. [106] optimize the Wasserstein distance between two sets of data points, one with labels and one without, with an additional penalty on the transport plan preventing two differently labeled points being sent to the same target. They also combine this with classification-based losses on the points with known labels.
@@
Thus, in the absence of closed form solutions, most works turn towards iterative descent algorithms to solve the optimization problem. Here, the most fundamental approach is to apply gradient descent or its stochastic extension which was used in [15, 41, 73, 74, 106, 213, 166, 105, 202, 212, 143, 141]. To speed-up the convergence of the descent, some works, such as [9, 37, 38, 103, 124, 183, 160, 161, 159, 162, 29], use acceleration techniques such as Nesterov’s acceleration or the Adam algorithm. Others [87, 168] use L-BFGS to approximate the Hessian of the objective and achieve faster convergence.
These iterative methods typically face two challenges: (i) computing the gradient, and (ii) ensuring that the learned metric is positive definite for explicit parametrizations. The first challenge traditionally required manual computation of the gradient of the objective which has been mostly replaced by the usage of automatic differentiation frameworks in more recent approaches. Furthermore, the objective might not admit a classical gradient, e.g. if it involves solving an optimization problem as in the case of optimal transport-based objectives. Thus, some works explicitly use subgradient methods [48, 105, 201, 143]. The second challenge is often solved via the parametrization of the metric (see Section˜4.1) and thus mostly comes up in the optimization for constant metrics. In these cases, it is typically [73, 202] solved by projected gradient descent, i.e. alternating a gradient step with projecting the metric matrix onto cone of positive definite matrices.